A compound of carbon,hydrogen and nitrogen contains the elements in the ratio of 18:2:7.Its empirical formula will be

Grade:12

2 Answers

Ramesh V
14 years ago

Let the compound be CaHbNc

and stomic wts. are 12,1,14 respectiveky for C,H,N

12a/(12a+b+14c) = 18/27   ... (1)

b/(12a+b+14c) = 2/27        ... (2)

14c/(12a+b+14c) = 14/27   ... (3)

on dividing 1 and 3 , we have ... a = 3c

on dividing 2 and 3 , we have ... b = 4c

so, the empirical formula is C3H4N

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult
to understand - post it here and we will get you the answer and detailed solution very quickly.We are all
IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Ramesh  
IIT Kgp - 05 batch





Pratham Ashish
14 years ago

hi,

elements

ratio

P  =ratio/corresponding atomic wt

P divided by least no. in previous column

Ratio of atoms in empirical formula

C

18

18/12 = 1.5

1.5/0.5 =3

3

H

2

2/1 = 2

2/ 0.5 = 4

4

N

7

7 / 14 = 0.5

.5 /.5 = 1

1

  so , empirical formula of compound would be =

                                                                 = C3 H4 N